# Searching Meta Reasoning Skeleton to Guide LLM Reasoning

### v1

@@ 1 Introduction
Large language model (LLM) demonstrate superior performance on complex tasks such as math Q&A when equipped with step-by-step reasoning ability (Wei et al., 2022; OpenAI, 2024; DeepSeek-AI, 2025). Researches on cognition divide reasoning into two levels: base reasoning (reasoning for problem directly) and meta reasoning (higher-level reasoning about how to reason) (Flavell, 1979). Meta reasoning, considered a unique ability of human cognition (Ackerman & Thompson, 2017), entails awareness of one’s reasoning process and the deliberate selection of reasoning strategies. For instance, when encountering difficulty with math problem, humans shift solution by thinking “This approach is not working; I should try another method…” or they may verify their reasoning steps by reflecting “Some steps may have errors. Let me check a previous step…” These behaviors do not directly solve the problem itself but instead organized as skeleton to guide the reasoning process.
Inspired by such human behaviors, previous studies proposed to incorporate meta reasoning into LLM to guide their reasoning process and thereby enhance performance on complex reasoning tasks (Gao et al., 2024; Qi et al., 2025; Sui et al., 2025; Liu et al., 2025). Recent approaches typically predefine a set of meta reasoning strategies for intermediate reasoning steps and employ manually designed structures (e.g. sequential, parallel and tree) to organize the strategies into meta reasoning skeleton. For example, rStar (Qi et al., 2025) and Meta-Reasoner (Sui et al., 2025) both define step-wise strategies such as decomposing question into sub-questions. rStar leverages Monte Carlo Tree Search (MCTS) (Coulom, 2006) to select and organize strategies, whereas Meta-Reasoner arranges them in a sequential way and selects at each step via multi-armed bandit (Gittins, 1979). An intuitive illustration of these manually designed skeleton is provided in Figure 2.
The aforementioned methods based on manually designed meta reasoning skeleton improved LLM reasoning performance. However, evidence from cognition science suggests that meta reasoning skeletons should vary for different queries, due to reasoner ability, query difficulty, discipline characteristic, etc. (Scott & Berman, 2013; Erickson & Heit, 2015; Rouault et al., 2018). For example in Figure 1, knowledge-intensive problems (Q3 about biology) rely more heavily on knowledge-recall strategy while shallower thinking depth than thinking-intensive problem (Q1 and Q2 about math).
@@
More difficult problems (Q1) may demand more parallel reasoning branches with solution exploration strategy than easier one (Q2). Besides, the logical dependency of reasoning steps can be too intricate (Besta et al., 2024) to capture by sequential, parallel, or tree-structured skeletons in prior works. The skeleton of Q1 involves parallel branches (steps 1–3 forming one branch while steps 4–6 another) and multiple dependency (step 6 simultaneously depends on step 5 as well as step 3 from early branches). Skeleton of Q3 summarizes two steps (step 2 and 3) to make it answer confident. The query-specific requirement and the intricate logical dependency among reasoning steps make it challenging for existing methods with limited manually designed meta reasoning skeletons (Figure 2) to work well across all queries.
Automated machine learning (AutoML) seeks to generate machine learning configurations for given task in a data-driven manner (Shen et al., 2024), thereby reducing the need for manual design and tuning for neural architectures (Elsken et al., 2019) and hyperparameter (Feurer & Hutter, 2019). Inspired by success of AutoML, we propose AutoMR, a framework that automatically searches for query-aware meta reasoning skeletons to guide LLM to reason for correct answer, where we represent meta reasoning skeleton as single-source edge-heterogeneous directed acyclic graph (DAG) to cover skeleton in prior works and capture intricate logical dependencies. Specifically, we first design an extensive DAG-based skeleton search space. Then we formulate the meta reasoning skeleton search problem, which poses two technical difficulties specific to query-aware skeleton search. The first is to derive any skeleton for given query from the extensive search space efficiently. The other is to adapt derived skeleton to evolving base reasoning context, considering inherent step-by-step property of reasoning process. To tackle the difficulties, we design a skeleton sampling algorithm that expands meta reasoning skeleton node by node dynamically based on base reasoning context at inference time. We prove that this algorithm introduces minimal additional computation overhead compared with naive LLM reasoning process. Compared with prior meta reasoning method, our search for meta reasoning skeleton improves reasoning performance. Moreover, we show that our search and inference algorithm is efficient theoretically and empirically.
@@ 2 Related Works
Meta Reasoning in LLM. Meta reasoning is an ability of human cognition involving determining reasoning strategy about how to reason (Flavell, 1979; Ackerman & Thompson, 2017). Previous works explored to introduce meta reasoning into LLM to guide it reasoning (Liu et al., 2025; Alazraki & Rei, 2025; Yan et al., 2025; Xiang et al., 2025; De Sabbata et al., 2024; Wan et al., 2025; Didolkar et al., 2024). Meta Reasoning Prompt (MRP) (Gao et al., 2024) includes classic strategies like CoT (Wei et al., 2022), Self-Refine (Madaan et al., 2023), etc. It first prompts LLM to choose one strategy for given query and then reason guided by that strategy. Strategies in MRP are holistic, meaning that MRP uses only one strategy for the whole reasoning process without adjusting when reasoning progressing. In contrast, recent methods usually use step-wise meta reasoning strategies (Yang et al., 2025a; b) and choose strategy for each step during reasoning. For example, rStar (Qi et al., 2025) define step-wise reasoning strategies such as proposing a sub-question, and then use MCTS to build tree-structured meta reasoning skeleton. Meta-Reasoner (Sui et al., 2025) also uses step-wise reasoning strategies but organizes them with sequential skeleton and uses multi-armed bandit to select strategy for each step. This kind of methods incorporate more fine-grained meta reasoning guidance and allow adjusting strategies during reasoning, thus performing better empirically than MRP.
Automated Machine Learning (AutoML). AutoML aims to search for high-performing machine learning (ML) configuration for given task automatically, reducing demand for human manual design (He et al., 2021) to adapt to task-specific requirement. Typical AutoML atomizes ML configurations to construct search space and develop search algorithm to find effective candidates (Shen et al., 2024). Previous works implemented this idea for multiple ML configurations such as neural architecture search (NAS) (White et al., 2023; Liu et al., 2019; Pham et al., 2018) and hyperparameter search (Yang & Shami, 2020; Shen et al., 2023), and have achieved success. For example, architectures found by NAS surpass human-designed ones on various tasks, such as computer vision (Real et al., 2019) and natural language processing (So et al., 2019). Recent works explored integrating AutoML with LLMs, like automating LLM agent workflow building (Zhuge et al., 2024; Zhang et al., 2025a; Saad-Falcon et al., 2025). However, applying AutoML method to search for meta reasoning skeleton is non-trivial due to factors specific to LLM reasoning task, including query-specific requirement, intricate logical dependency, and evolving reasoning context.
@@ 3.1 Search Space
Based on this unified view, we construct search space to contain all skeletons represented by single-source edge-heterogeneous DAG as shown in Figure 2, as long as the sum of tokens for all node content except source node (i.e. number of tokens generate by LLM) dose not reach token budget $\mathcal{B}$}, where $\mathcal{B}$} is a hyperparameter.
We summarize the meta reasoning behaviors in previous works about LLM reasoning (Gandhi et al., 2025; Chen et al., 2025), which gives meta reasoning strategy set ${\mathcal{S}}=\{\textbf{Next},\textbf{Reflect},\textbf{Explore},\textbf{Decompose},\textbf{Summarize},\textbf{Recall},\textbf{Answer}\}$. All of these meta reasoning strategies are implemented by designed prompt. Functions and prompt of these strategies are summarized in Table 2 in Appendix A.1. Following previous works (Liu et al., 2019), we also introduce a special zero edge type to indicate an edge in fact dose not exists.
@@ 3.2 Search Strategy
When implementing a policy $P$ for deriving a query-aware skeleton, this search problem poses two technical challenges specific to LLM reasoning. First, the search space is extensive, so the derivation procedure must efficiently explore it to recover arbitrary skeletons in it. Second, because reasoning process unfolds step by step (Wei et al., 2022; Nye et al., 2021), the derivation process should adapt meta reasoning strategy at each step in skeleton to evolving base reasoning context, rather than fixing the skeleton a priori before reasoning for given query.
@@ 3.2.2 Overall Search Algorithm
With $P_{\theta}(\alpha_{q}|q)$ defined in (3), we follow REINFORCE (Williams, 1992; Zoph & Le, 2017), a policy gradient algorithm implementing unbiased empirical approximation of objective, to optimize $\theta$. Specifically, we sample batches with $N$ query-answer pairs $(q_{i},a_{i})$ from training set each time and optimize $\theta$ with these batches iteratively. For each $(q_{i},a_{i})$ in batch, we sample $M$ skeletons $\alpha_{q_{i}}^{j}$ from $P_{\theta}(\cdot|q_{i})$ and evaluate their performance with $r(\cdot)$ respectively. The update to $\theta$ in each iteration by estimated policy gradient with a batch is as follows, where $\eta$ is learning rate:
@@ 3.3 Technical Comparison with AutoML
Different from prior meta reasoning methods that rely on manually designed skeleton (Qi et al., 2025; Sui et al., 2025), AutoMR draws inspiration from AutoML to search for query-aware meta reasoning skeleton from DAG-based search space, thereby addressing query-specific requirements. Technically, AutoMR is related to topics in AutoML such as neural architecture search. Recent studies have extended AutoML ideas to LLM-related tasks, such as automating agent workflow building (Zhuge et al., 2024; Zhang et al., 2025a). However, the unique properties of LLM reasoning tasks make AutoMR particularly suited for meta reasoning skeleton search. First, reasoning queries often exhibit highly specific demands, making a single meta reasoning skeleton insufficient. Second, the reasoning process typically involves intricate logical dependencies. Third, reasoning unfolds step by step, with the base reasoning context dynamically evolving as each new step is generated. These characteristic fundamentally differs from those of neural architecture or agent workflow, which is usually fixed for all queries or static during inference. For example, Prior approaches (Zoph et al., 2018; Zhuge et al., 2024) generally output a single architecture or agent workflow for all queries. While instance-aware methods (Cheng et al., 2020; Zhang et al., 2025a) produce input-specific architecture or workflow that remain static during inference. Such differences in task properties makes the search techniques in these methods perform well in their target scenarios but cannot be applied to meta reasoning skeleton search directly. We compare these search techniques empirically by ablation study in Section 4.3.
@@ 4.1 Setup
Baselines. We implement the following types of baselines: (1) Classic methods, including Direct-I/O and CoT (Wei et al., 2022). (2) Meta reasoning methods, including MRP (Gao et al., 2024), rStar (Qi et al., 2025) and Meta-Reasoner (Sui et al., 2025). We also include MaAS (Zhang et al., 2025a), a method using NAS technique to automate multi-agent workflow building.
AutoMR and all the baselines are implemented based on two LLMs including LLaMA3.2-3B-Inst (hereinafter referred to as ”LLaMA”) (Meta-AI, 2024) and Qwen2.5-3B-Inst (hereinafter referred to as ”Qwen”) (Qwen-Team, 2025) to avoid impact on experimental results caused by unique properties of specific LLM (Gandhi et al., 2025). We set the same token budget to 1024 for all methods to ensure fair comparison. More implementation details of the baselines are introduced in Appendix C.1.
Datasets and Metric. We evaluate AutoMR and baselines on two domains, i.e. math Q&A and general multiple-choice. For math Q&A, we choose GSM8K (Cobbe et al., 2021), MATH-500 (Hendrycks et al., 2021), AMC (including AMC 2022 and AMC 2023) and Olympiad (only open-ended text-only math subset to avoid influence from multi-modal and multilingual input information) (He et al., 2024) to evaluate. We use training split of MATH dataset to train AutoMR and baselines that need training. For general multiple-choice, we choose MMLU-Pro (Wang et al., 2024) and split it into four subsets as Science, Humanities, Social and Other referring to Zhang et al. (2025b), to evaluate. We collect training split of MMLU-Pro to train. Details of these datasets are summarized in Appendix C.2. We use Accuracy as metric to evaluate these methods .
@@ 4.2 Performance Comparison
Table 1: The overall performance on math Q&A an general multi-choice. Letters after method names means the used skeleton structure. S: Sequential; T: Tree; G: DAG; “-” means not applicable.
@@ 4.3 Ablation Study
Influence of token budget scaling. Previous works shows that LLM reasoning performance improves whentoken budget increases (OpenAI, 2024; Snell et al., 2025). We evaluate the performance when scaling token budget $\mathcal{B}$}. We compare AutoMR with baselines able to scale token budget. Specifically, for CoT we implement sequential scaling technique Budget Forcing (Muennighoff et al., 2025) and parallel technique Majority Voting (Wang et al., 2023). We also choose Meta-Reasoner and rStar as baselines. We do not include MaAS as baselines to evaluate because it do not provide scaling technique in original paper. The scaling technique implementation details of these methods are in Appendix C.1. We evaluate on MATH-500 and Science based on Qwen. According to results in Figure 3, we observe that when token budget increases, each method improves performance on the whole. Specifically, the scaling efficiency on knowledge-intensive Science subet is much slower than that on thinking-intensive MATH-500, according with recent research (Zhao et al., 2025). Forcing sequential scaling (i.e. Budget Forcing and Meta-Reasoner) scale slowly. Majority Voting based on parallel skeleton and rStar based on tree-structured skeleton scale more efficiently than sequential ones. AutoMR achieve the highest scaling efficiency, because search space based on DAG in AutoMR allows more extensive skeleton exploration.
@@
Effectiveness of search strategy. We evaluate the effectiveness of search strategy in Section 3.2 against Random Search (RS(Bergstra & Bengio, 2012), a common AutoML baseline (Li & Talwalkar, 2020). We also assess effectiveness of dynamic skeleton sampling algorithm by comparing it with two variants. Query-Invariant (QI), sampling single meta reasoning skeleton shared by all queries of a task, as in prior NAS methods (Liu et al., 2019; Pham et al., 2018). Complete in Advance (CA), sampling query-specific skeletons before reasoning starts but not based on reasoning context (Cheng et al., 2020; Zhang et al., 2025a). Implementation details of these sampling methods are in Appendix C.1. We compare them on MATH-500 and Science.
Figure 5: Ablation study on search strategy.
@@
Training and inference efficiency. To support theoretical analysis in Section 3.2.2 that AutoMR incurs minimal additional computation, we evaluate both training and inference costs of AutoMR and baselines requiring training, including Meta-Reasoner and MaAS, based on both Qwen and LLaMA on MATH-500 dataset. We also implement GRPO (Shao et al., 2024), a reinforcement learning method to to enhance LLM reasoning, based on LoRA (Hu et al., 2022) as a baseline in our experiment setting. Results in Figure 4 show training cost (x-axis), performance on MATH-500 (y-axis), and inference cost (circle area). In terms of training, AutoMR and other two baselines require far less time than GRPO, which fine-tunes LLM parameters directly. However, only AutoMR achieves comparable performance with Qwen and even surpasses it with LLaMA. In terms of inference, AutoMR is slightly slower than naive reasoning process based on GRPO-trained LLM and slightly faster than MaAS, while being substantially more efficient than Meta-Reasoner, which relies on additional LLM calls to summarize reasoning progress. Instead, AutoMR employs a lightweight MLP to process representations produced during reasoning, avoiding extra LLM calls.

### v2

@@ 1 Introduction
Large language model (LLM) demonstrate superior performance on complex tasks such as math Q&A when equipped with step-by-step reasoning ability (cot; o1; r1). Researches on cognition divide reasoning into two levels: base reasoning (reasoning for problem directly) and meta reasoning (higher-level reasoning about how to reason) (metacognition). Meta reasoning, considered a unique ability of human cognition (ackerman2017meta), entails awareness of one’s reasoning process and the deliberate selection of reasoning strategies. For instance, when encountering difficulty with math problem, humans shift solution by thinking “This approach is not working; I should try another method…” or they may verify their reasoning steps by reflecting “Some steps may have errors. Let me check a previous step…” These behaviors do not directly solve the problem itself but instead organized as skeleton to guide the reasoning process.
Inspired by such human behaviors, previous studies proposed to incorporate meta reasoning into LLM to guide their reasoning process and thereby enhance performance on complex reasoning tasks (mrp; rstar; metareasoner; metascale). Recent approaches typically predefine a set of meta reasoning strategies for intermediate reasoning steps and employ manually designed structures (e.g. sequential, parallel and tree) to organize the strategies into meta reasoning skeleton. For example, rStar (rstar) and Meta-Reasoner (metareasoner) both define step-wise strategies such as decomposing question into sub-questions.
@@
rStar leverages Monte Carlo Tree Search (MCTS) (mcts) to select and organize strategies, whereas Meta-Reasoner arranges them in a sequential way and selects at each step via multi-armed bandit (mab). An intuitive illustration of these manually designed skeleton is provided in Figure 2.
The aforementioned methods based on manually designed meta reasoning skeleton improved LLM reasoning performance. However, evidence from cognition science suggests that meta reasoning skeletons should vary for different queries, due to reasoner ability, query difficulty, discipline characteristic, etc. (scott2013examining; erickson2015metacognition; rouault2018human). For example in Figure 1, knowledge-intensive problems (Q3 about biology) rely more heavily on knowledge-recall strategy while shallower thinking depth than thinking-intensive problem (Q1 and Q2 about math). More difficult problems (Q1) may demand more parallel reasoning branches with solution exploration strategy than easier one (Q2). Besides, the logical dependency of reasoning steps can be too intricate (besta2024graph) to capture by sequential, parallel, or tree-structured skeletons in prior works. The skeleton of Q1 involves parallel branches (steps 1–3 forming one branch while steps 4–6 another) and multiple dependency (step 6 simultaneously depends on step 5 as well as step 3 from early branches). Skeleton of Q3 summarizes two steps (step 2 and 3) to make it answer confident. The query-specific requirement and the intricate logical dependency among reasoning steps make it challenging for existing methods with limited manually designed meta reasoning skeletons (Figure 2) to work well across all queries.
Automated machine learning (AutoML) seeks to generate machine learning configurations for given task in a data-driven manner (automl), thereby reducing the need for manual design and tuning for neural architectures (nas) and hyperparameter (feurer2019hyperparameter). Inspired by success of AutoML, we propose AutoMR, a framework that automatically searches for query-aware meta reasoning skeletons to guide LLM to reason for correct answer, where we represent meta reasoning skeleton as single-source edge-heterogeneous directed acyclic graph (DAG) to cover skeleton in prior works and capture intricate logical dependencies. Specifically, we first design an extensive DAG-based skeleton search space. Then we formulate the meta reasoning skeleton search problem, which poses two technical difficulties specific to query-aware skeleton search. The first is to derive any skeleton for given query from the extensive search space efficiently. The other is to adapt derived skeleton to evolving base reasoning context, considering inherent step-by-step property of reasoning process. To tackle the difficulties, we design a skeleton sampling algorithm that expands meta reasoning skeleton node by node dynamically based on base reasoning context at inference time. We prove that this algorithm introduces minimal additional computation overhead compared with naive LLM reasoning process. Compared with prior meta reasoning method, our search for meta reasoning skeleton improves reasoning performance. Moreover, we show that our search and inference algorithm is efficient theoretically and empirically.
@@ 2 Related Works
Meta Reasoning in LLM. Meta reasoning is an ability of human cognition involving determining reasoning strategy about how to reason (metacognition; ackerman2017meta). Previous works explored to introduce meta reasoning into LLM to guide it reasoning (metascale; alazraki-rei-2025-meta; yan2025position; metacot; de2024rational; rema; didolkar2024metacognitive). Meta Reasoning Prompt (MRP) (mrp) includes classic strategies like CoT (cot), Self-Refine (selfrefine), etc. It first prompts LLM to choose one strategy for given query and then reason guided by that strategy. Strategies in MRP are holistic, meaning that MRP uses only one strategy for the whole reasoning process without adjusting when reasoning progressing. In contrast, recent methods usually use step-wise meta reasoning strategies (reasonflux; yang2025supercorrect) and choose strategy for each step during reasoning. For example, rStar (rstar) define step-wise reasoning strategies such as proposing a sub-question, and then use MCTS to build tree-structured meta reasoning skeleton. Meta-Reasoner (metareasoner) also uses step-wise reasoning strategies but organizes them with sequential skeleton and uses multi-armed bandit to select strategy for each step. This kind of methods incorporate more fine-grained meta reasoning guidance and allow adjusting strategies during reasoning, thus performing better empirically than MRP.
Automated Machine Learning (AutoML). AutoML aims to search for high-performing machine learning (ML) configuration for given task automatically, reducing demand for human manual design (he2021automl) to adapt to task-specific requirement. Typical AutoML atomizes ML configurations to construct search space and develop search algorithm to find effective candidates (automl). Previous works implemented this idea for multiple ML configurations such as neural architecture search (NAS) (nas1000; darts; enas) and hyperparameter search (yang2020hyperparameter; shen2023efficient), and have achieved success. For example, architectures found by NAS surpass human-designed ones on various tasks, such as computer vision (real2019regularized) and natural language processing (so2019evolved). Recent works explored integrating AutoML with LLMs, like automating LLM agent workflow building (zhuge2024gptswarm; zhang2025multiagent; saad-falcon2025an). However, applying AutoML method to search for meta reasoning skeleton is non-trivial due to factors specific to LLM reasoning task, including query-specific requirement, intricate logical dependency, and evolving reasoning context.
@@ 3.1 Search Space
Based on this unified view, we construct search space to contain all skeletons represented by single-source edge-heterogeneous DAG as shown in Figure 2, as long as the sum of tokens for all node content except source node (i.e. number of tokens generate by LLM) does not reach token budget $\mathcal{B}$}, where $\mathcal{B}$} is a hyperparameter.
We summarize the meta reasoning behaviors in previous works about LLM reasoning (gandhi2025cognitivebehaviorsenableselfimproving; chen2025reasoningerasurveylong), which gives meta reasoning strategy set ${\mathcal{S}}=\{\textbf{Next},\textbf{Reflect},\textbf{Explore},\textbf{Decompose},\textbf{Summarize},\textbf{Recall},\textbf{Answer}\}$. All of these meta reasoning strategies are implemented by designed prompt. Functions and prompt of these strategies are summarized in Table 3 in Appendix A.1. Following previous works (darts), we also introduce a special zero edge type to indicate an edge in fact does not exists.
@@ 3.2 Search Strategy
When implementing a policy $P$ for deriving a query-aware skeleton, this search problem poses two technical challenges specific to LLM reasoning. First, the search space is extensive, so the derivation procedure must efficiently explore it to recover arbitrary skeletons in it. Second, because reasoning process unfolds step by step (cot; nye2022show), the derivation process should adapt meta reasoning strategy at each step in skeleton to evolving base reasoning context, rather than fixing the skeleton a priori before reasoning for given query.
@@ 3.2.2 Overall Search Algorithm
With $P_{\theta}(\alpha_{q}|q)$ defined in (3), we follow REINFORCE (reinforce; zoph2017neural), a policy gradient algorithm implementing unbiased empirical approximation of objective, to optimize $\theta$. Specifically, we sample batches with $N$ query-answer pairs $(q_{i},a_{i})$ from training set each time and optimize $\theta$ with these batches iteratively. For each $(q_{i},a_{i})$ in batch, we sample $M$ skeletons $\alpha_{q_{i}}^{j}$ from $P_{\theta}(\cdot|q_{i})$ and evaluate their performance with $r(\cdot)$ respectively. The update to $\theta$ in each iteration by estimated policy gradient with a batch is as follows, where $\eta$ is learning rate:
@@ 3.3 Technical Comparison with AutoML
Different from prior meta reasoning methods that rely on manually designed skeleton (rstar; metareasoner), AutoMR draws inspiration from AutoML to search for query-aware meta reasoning skeleton from DAG-based search space, thereby addressing query-specific requirements. Technically, AutoMR is related to topics in AutoML such as neural architecture search. Recent studies have extended AutoML ideas to LLM-related tasks, such as automating agent workflow building (zhuge2024gptswarm; zhang2025multiagent). However, the unique properties of LLM reasoning tasks make AutoMR particularly suited for meta reasoning skeleton search. First, reasoning queries often exhibit highly specific demands, making a single meta reasoning skeleton insufficient. Second, the reasoning process typically involves intricate logical dependencies. Third, reasoning unfolds step by step, with the base reasoning context dynamically evolving as each new step is generated. These characteristic fundamentally differs from those of neural architecture or agent workflow, which is usually fixed for all queries or static during inference. For example, Prior approaches (zoph2018learning; zhuge2024gptswarm) generally output a single architecture or agent workflow for all queries. While instance-aware methods (cheng2020instanas; zhang2025multiagent) produce input-specific architecture or workflow that remain static during inference. Such differences in task properties makes the search techniques in these methods perform well in their target scenarios but cannot be applied to meta reasoning skeleton search directly. We compare these search techniques empirically by ablation study in Section 4.3.
@@ 4.1 Setup
Baselines. We implement the following types of baselines: (1) Classic methods, including Direct-I/O and CoT (cot). (2) Meta reasoning methods, including MRP (mrp), rStar (rstar) and Meta-Reasoner (metareasoner). We also include MaAS (zhang2025multiagent), a method using NAS technique to automate multi-agent workflow building.
AutoMR and all the baselines are implemented based on two LLMs including LLaMA3.2-3B-Inst (hereinafter referred to as ”LLaMA”) (llama3herdmodels) and Qwen2.5-3B-Inst (hereinafter referred to as ”Qwen”) (qwen25) to avoid impact on experimental results caused by unique properties of specific LLM (gandhi2025cognitivebehaviorsenableselfimproving). We set the same token budget to 1024 for all methods to ensure fair comparison. More implementation details of the baselines are introduced in Appendix C.1.
Datasets and Metric. We evaluate AutoMR and baselines on two domains, i.e. math Q&A and general multiple-choice. For math Q&A, we choose GSM8K (cobbe2021gsm8k), MATH-500 (hendrycksmath2021), AMC (including AMC 2022 and AMC 2023) and Olympiad (only open-ended text-only math subset to avoid influence from multi-modal and multilingual input information) (he2024olympiadbench) to evaluate. We use training split of MATH dataset to train AutoMR and baselines that need training. For general multiple-choice, we choose MMLU-Pro (wang2024mmlupro) and split it into four subsets as Science, Humanities, Social and Other referring to zhang2025right, to evaluate. We collect training split of MMLU-Pro to train. Details of these datasets are summarized in Appendix C.2. We use Accuracy as metric to evaluate these methods .
@@ 4.2 Performance Comparison
Table 1: The overall performance on math Q&A and general multi-choice. Letters after method names means the used skeleton structure. S: Sequential; T: Tree; G: DAG; “-” means not applicable.
@@ 4.3 Ablation Study
Influence of token budget scaling. Previous works shows that LLM reasoning performance improves whentoken budget increases (o1; tts). We evaluate the performance when scaling token budget $\mathcal{B}$}. We compare AutoMR with baselines able to scale token budget. Specifically, for CoT we implement sequential scaling technique Budget Forcing (s1) and parallel technique Majority Voting (sc). We also choose Meta-Reasoner and rStar as baselines. We do not include MaAS as baselines to evaluate because it do not provide scaling technique in original paper. The scaling technique implementation details of these methods are in Appendix C.1. We evaluate on MATH-500 and Science based on Qwen. According to results in Figure 3, we observe that when token budget increases, each method improves performance on the whole. Specifically, the scaling efficiency on knowledge-intensive Science subset is much slower than that on thinking-intensive MATH-500, according with recent research (zhao2025test). Forcing sequential scaling (i.e. Budget Forcing and Meta-Reasoner) scale slowly. Majority Voting based on parallel skeleton and rStar based on tree-structured skeleton scale more efficiently than sequential ones. AutoMR achieve the highest scaling efficiency, because search space based on DAG in AutoMR allows more extensive skeleton exploration.
@@
Effectiveness of search strategy. We evaluate the effectiveness of search strategy in Section 3.2 against Random Search (RS(randomsearch), a common AutoML baseline (li2020random). We also assess effectiveness of dynamic skeleton sampling algorithm by comparing it with two variants. Query-Invariant (QI), sampling single meta reasoning skeleton shared by all queries of a task, as in prior NAS methods (darts; enas). Complete in Advance (CA), sampling query-specific skeletons before reasoning starts but not based on reasoning context (cheng2020instanas; zhang2025multiagent). Implementation details of these sampling methods are in Appendix C.1. We compare them on MATH-500 and Science.
Table 2: Ablation study on search strategy.
@@
Training and inference efficiency. To support theoretical analysis in Section 3.2.2 that AutoMR incurs minimal additional computation, we evaluate both training and inference costs of AutoMR and baselines requiring training, including Meta-Reasoner and MaAS, based on both Qwen and LLaMA on MATH-500 dataset. We also implement GRPO (deepseekmath), a reinforcement learning method to enhance LLM reasoning, based on LoRA (lora) as a baseline in our experiment setting. Results in Figure 4 show training cost (x-axis), performance on MATH-500 (y-axis), and inference cost (circle area). In terms of training, AutoMR and other two baselines require far less time than GRPO, which fine-tunes LLM parameters directly. However, only AutoMR achieves comparable performance with Qwen and even surpasses it with LLaMA. In terms of inference, AutoMR is slightly slower than naive reasoning process based on GRPO-trained LLM and slightly faster than MaAS, while being substantially more efficient than Meta-Reasoner, which relies on additional LLM calls to summarize reasoning progress. Instead, AutoMR employs a lightweight MLP to process representations produced during reasoning, avoiding extra LLM calls.